OpenStudy (anonymous):

f(x) = -3x^2 + 2x - 7 f(-1) =

OpenStudy (anonymous):

substitute -1 instead of x in the function and calculate

OpenStudy (anonymous):

oh!! okay i get it!

OpenStudy (anonymous):

f(-1)=-3*(-1)^2+2(-2)-7=-3*1-4-7=3-11=-8 the answer is -8

OpenStudy (anonymous):

thts not one of the answers! i did it and got -12 and thats one of the answers

OpenStudy (anonymous):

opps yeas. you are right @marie123 . I'm mistaken!!! sorry

OpenStudy (anonymous):

its okay! :) thank you regardless!

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

plug 1 into the fuction f(1)=-3(-1)^2 + 2(-1) - 7 =-3+(-2)-7 =-5-7 =-12